OpenStudy (anonymous):

if x-y=11, and x^2-y^2=22, what is 5x+5y? please explain step by step!! thanks!

OpenStudy (anonymous):

x - y = 11 x^2-y^2=(x+y) (x-y) = 11 (x+y) = 22 x+y = 2 5x + 5y = 10